A typical chemical formula for aerobic growth of a microorganisms is


C2H5OH + a(O2) + (b)NH3 —> (c)CH1.7H0.15O0.4 + (d)H2O + (e)CO2


where term CH1.7H0.15O0.4 represents metabolism of the microorganism. The ratio of moles of CO2 produced per mole of O2 consumed is called the respiratory quotient, RQ, which can be determined experimentally. Given this ratio we have 4 constants, a-d that are unknown. We can perform a mass balance on each of the four key elements

Carbon: 2 = c + (RQ)a

Hydrogen: 6 + 3b = 1.7c + 2d

Oxygen: 1 + 2a = 0.4c + d + 2(RQ)a

Nitrogen: b = 0.15c 

Let RQ = 0.8 and then find the vector [a b c d] using Gauss Elimination Method.



1
Expert's answer
2022-01-02T08:16:09-0500

The equations may be written as

"\\begin{matrix}\n\\text{Carbon: }&\t(RQ)a + c = 2 \\\\\n\\text{Hydrogen: }& -3b + 1.7c + 2d = 6\\\\\n\\text{Oxygen: }& (2(RQ)-2)a + 0.4c + d = 1\\\\\n\\text{Nitrogen: }& b - 0.15c = 0\n\\end{matrix}"


Or substituting RQ = 0.8 in matrix form


"\begin{bmatrix}
 0.8 & 0 & 1 & 0 \\
 0 & -3 & 1.7 & 2 \\
 -0.4 & 0 & 0.4 & 1 \\
 0 & 1 & -0.15 & 0
\end{bmatrix} 
\begin{bmatrix}
 a \\
 b \\
 c \\
 d 
\end{bmatrix} =
\begin{bmatrix}
 2 \\
 6 \\
 1 \\
 0
\end{bmatrix}"


Adding 1/2 of the first row to the third one:


"\begin{bmatrix}
 0.8 & 0 & 1 & 0 \\
 0 & -3 & 1.7 & 2 \\
 0 & 0 & 0.9 & 1 \\
 0 & 1 & -0.15 & 0
\end{bmatrix} 
\begin{bmatrix}
 a \\
 b \\
 c \\
 d 
\end{bmatrix} =
\begin{bmatrix}
 2 \\
 6 \\
 2 \\
 0
\end{bmatrix}"


Add 1/3 of the second row to the fourth:


"\begin{bmatrix}
 0.8 & 0 & 1 & 0 \\
 0 & -3 & 1.7 & 2\\
 0 & 0 & 0.9 & 1 \\
 0 & 0 & 0.417 & 0.667
\end{bmatrix} 
\begin{bmatrix}
 a \\
 b \\
 c \\
 d 
\end{bmatrix} =
\begin{bmatrix}
 2 \\
 6 \\
 2 \\
 2
\end{bmatrix}"


Subtract 4.17/9 of the third row from the fourth:


"\begin{bmatrix}
 0.8 & 0 & 1 & 0 \\
 0 & -3 & 1.7 & 2\\
 0 & 0 & 0.9 & 1 \\
 0 & 0 & 0 & 0.204
\end{bmatrix} 
\begin{bmatrix}
 a \\
 b \\
 c \\
 d 
\end{bmatrix} =
\begin{bmatrix}
 2 \\
 6 \\
 2 \\
 1.07
\end{bmatrix}"


Therefor


"d = 1.07 / 0.204 \\
c = (2 - d) / 0.9 \\
b = (6 - 1.7 c - 2 d) / (-3) \\
a = (2 - c) / 0.8"


"\begin{bmatrix}
 a \\
 b \\
 c \\
 d
\end{bmatrix} =
\begin{bmatrix}
 7.0 \\
 -0.55 \\
 -3.6 \\
5.24
\end{bmatrix}"